calculate the percent by mass of solute at 8.60 g of salt and 95.0 g of solution
1
Expert's answer
2021-01-06T05:54:58-0500

%(m/m) = "\frac{mass of solute }{mass of solution}" × 100%


= "\frac{8.60g}{95.0g}" × 100%


= 9.052 %